two gases No and oxygen were introduced at the both ends of one meter tube.At what distance from No gas end ,brown gas will be seen?

Grade:

5 Answers

Sudheesh Singanamalla
13 years ago

Nitrogen di oxide (No)  ; i didn't understand what No is but i think it is Nitrogen Oxide right ? , so assuming that it is NO , i m continuing with the solution.

in a one metre tube , Nitrogen Di Oxide and Oxygen is passed through each side.

2NO + O2 → 2NO2 (because NO2 Is the brown gas)

therefore by the reaction stoichiometry we see that

2 moles of NO are required for every 1 mole of O to for NO2

SO ratio of NO to total = 2 : 3

the brown gas will be found near the part of the metre tube where oxygen is being pumped from.

it will be seen at a distance of 2/3 * 100 = 66.667 cm from the point where Nitrogen di oxide is introduced.

so it is 33.333cm from the end where oxygen is introduced !

 

Please approve !

Sudheesh Singanamalla
13 years ago

Don't post the same questions too many times

Sudheesh Singanamalla
13 years ago

Your question has already been answered

vikas askiitian expert
13 years ago

rate of diffusion = k/(molar mass)1/2

r02 = k/root16

rNO = k/root14

rNO/rO2 = root(8/7)

xNO/XO2 = root8/7

xNO + xO2 = 1

XNo = 1/1+root8/7

  =  0.483m

Anmol Vijaywargiya
13 years ago

Let the distance be x metres from NO2.... as x is directly proportional to 1/√Molecular weight x/1-x =√M(O2)/√M(NO2)